Solve the equation x^3+2x^2+x+2=0, if -2 is a root.
Answer by checkley77(12844)

x^3+2x^2+x+2=0, if -2 is a root.
-2^3+2*-2^2-2+2=0
-8+2*+4-2+2=0
-8+8-2+2=0
0+0=0
0=0